# Flat band driven itinerant magnetism in the Co-pnictides (La,Ca)Co2(As,P)2

### v1

@@ II Acknowledgements
D.S., A.K. and S.B-C. acknowledge financial support from the MINECO of Spain through the project PID2021-122609NB-C21 and by MCIN and by the European Union Next Generation EU/PRTR-C17.I1, as well as by IKUR Strategy under the collaboration agreement between IKERBASQUE Foundation and DIPC on behalf of the Department of Education of the Basque Government. C-Y.L. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 101020833). M.G.D. acknowledges financial support from Government of the Basque Country through the pre-doctoral fellowship PRE 2023 2 0024. M.S. acknowledges the support from the National Science Foundation (award DMR-2233902). This research used resources of the ESM beamline of the National Synchrotron Light Source II, a U.S. Department of Energy (DOE) Office of Science User Facility operated for the DOE Office of Science by Brookhaven National Laboratory under Contract No. DE-SC0012704. LOREA beamline is co-funded by the European Regional Development Fund (ERDF) within the ”Framework of the Smart Growth Operative Programme 2014-2020”. We acknowledge the MAX IV Laboratory for beamtime on the Bloch beamline under proposal 20210823. Research conducted at MAX IV, a Swedish national user facility, is supported by Vetenskapsrådet (Swedish Research Council, VR) under contract 2018-07152, Vinnova (Swedish Governmental Agency for Innovation Systems) under contract 2018-04969 and Formas under contract 2019-02496. P.G. acknowledges financial support from the MINECO project PID2023-149494NB-C32. M.G.V. thanks support to the Spanish Ministerio de Ciencia e Innovación (PID2022-142008NB-I00), the Canada Excellence Research Chairs Program for Topological Quantum Matter and the Diputación Foral de Gipuzkoa Programa Mujeres y Ciencia.

### v2

@@ Acknowledgments
D.S., A.K. and S.B-C. acknowledge financial support from the MINECO of Spain through the project PID2021-122609NB-C21 and by MCIN and by the European Union Next Generation EU/PRTR-C17.I1, as well as by IKUR Strategy under the collaboration agreement between IKERBASQUE Foundation and DIPC on behalf of the Department of Education of the Basque Government. C-Y.L. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 101020833). M.G.D. acknowledges financial support from the Government of the Basque Country through the pre-doctoral fellowship PRE 2023 2 0024. M.S. acknowledges the support from the National Science Foundation (award DMR-2233902). This research used resources of the ESM beamline of the National Synchrotron Light Source II, a U.S. Department of Energy (DOE) Office of Science User Facility operated for the DOE Office of Science by Brookhaven National Laboratory under Contract No. DE-SC0012704. LOREA beamline is co-funded by the European Regional Development Fund (ERDF) within the ”Framework of the Smart Growth Operative Programme 2014-2020”. We acknowledge the MAX IV Laboratory for beamtime on the Bloch beamline under proposal 20210823. Research conducted at MAX IV, a Swedish national user facility, is supported by Vetenskapsrådet (Swedish Research Council, VR) under contract 2018-07152, Vinnova (Swedish Governmental Agency for Innovation Systems) under contract 2018-04969 and Formas under contract 2019-02496. P.G. acknowledges financial support from the MINECO project PID2023-149494NB-C32. M.G.V. thanks support to the Spanish Ministerio de Ciencia e Innovación (PID2022-142008NB-I00), the Canada Excellence Research Chairs Program for Topological Quantum Matter, and the Diputación Foral de Gipuzkoa Programa Mujeres y Ciencia.